# Catalog of outbursts of neutron star LMXBs

### v1

@@ 3.1 General trends
In Fig. 1 we show the peak fluxes from all transient neutron star LMXBs (a few lie outside our bounds), with the most frequent outbursters indicated with different colors. In Fig. 2, we show the number of outbursts per year. We first review key information about the relevant monitoring instruments.
RXTE/ASM monitoring produced a dramatic increase in detected outbursts, from January 1996 until 2009-2011  [^7] The RXTE/ASM sensitivity decayed from 2009 until RXTE was turned off in 2011.. RXTE/ASM achieved roughly 15 mCrab ($F_{X}=3\times 10^{-10}$ ergs/cm2/s) 3-sigma sensitivity per day, in 2-10 keV (Levine et al., 1996). Indeed, detections from 1996 on now did reach $3\times 10^{-10}$, though completeness may have been a little higher, perhaps $6\times 10^{-10}$ ergs/cm2/s. The RXTE/ASM covered most sources daily, except for about 30 days each year when the Sun was nearest for sources in the ecliptic.
From 1996 to 2002, in each spring and fall, BeppoSAX conducted regular monitoring of the Galactic Bulge with its Wide-Field Camera (WFC), with a 40o by 40o field of view, 5’ angular resolution, and daily sensitivity of about 15 mCrab in 2-28 keV (Heise et al., 1999; in’t Zand et al., 2004). BeppoSAX’s WFC discovered a large number of faint ($L_{X}<10^{37}$ erg/s) X-ray transients in the Galactic Bulge, and also several X-ray bursts from transients that were too faint for the WFC to detect individually (Cornelisse et al., 2002b).
From 1999 to 2011, RXTE also conducted monitoring of the Galactic Bulge using slews of the PCA instrument (Swank & Markwardt, 2001). In practice, these typically reached a sensitivity of 5 mCrab, $10^{-10}$ ergs/cm2/s, 2.5-10 keV. Notably, the RXTE/PCA was able to detect many faint outbursts of NGC 6440 X-2 in 2009-2011.
The INTEGRAL Galactic Bulge hard X-ray monitoring program was particularly productive, with 12’ angular resolution and 5-15 mCrab daily sensitivity in 20-60 keV (Kuulkers et al., 2007), though it could observe the Bulge only 40% of the year, in the spring and fall. Twelve of these transients have principal names starting with IGR (only XTE has more, with 13), suggesting INTEGRAL’s impact. There appears (Fig. 2) to be a small decline in detections in the past few years, which may be due to declining sensitivity of INTEGRAL’s IBIS/ISGRI instrument. [^8] The Integral Galactic Bulge Monitoring Program’s website, http://integral.esac.esa.int/BULGE/, presents 18-40 keV IBIS/ISGRI lightcurves for sources, which show increasing noise from 2016 and stop in 2018.
Swift/BAT’s hard X-ray (15-50 keV) monitoring program has operated since 2006, with a daily 16 mCrab detection limit (Krimm et al., 2013). MAXI/GSC’s 2-20 keV monitor, with a 9 mCrab daily 3-sigma sensitivity (Sugizaki et al., 2013), has been operating since 2009.
@@
From Fig. 2, we see that 1996 marks a dramatic change in detected outburst frequency. A significant peak appears around 2009-2011 in all outbursts. Fig. 2 of Tetarenko et al. (2016b), transient outbursts of BH LMXBs, reveals a similar peak in 2007-2012. However, this peak is not significant in brighter NS LMXB outbursts with $F_{X}>3\times 10^{-10}$ ergs/s/cm2, suggesting that fainter outbursts were more detectable in that time frame (e.g. by the RXTE PCA bulge monitoring program; which detected, e.g., the NGC 6440 X-2 outbursts at $1-2\times 10^{-10}$ ergs/s/cm2; Heinke et al. 2010).
@@
We then construct a histogram of fluxes from all outbursts since 1996 (Fig. 3, blue triangles). This is sharply peaked at $F_{X}=10^{-9}$ ergs/s/cm2, which might suggest that our data becomes highly incomplete below $F_{X}=10^{-9}$ ergs/s/cm2. However, if we exclude the four most frequent outbursters (Aquila X-1, 4U 1608-52, MXB 1730-335, and GRS 1747-312), all of which have relatively long orbital periods, high mass transfer rates, and high outburst fluxes, we see a rather different histogram (Fig. 3, red squares), which does not peak until the bin $F_{X}=2-4\times 10^{-10}$ ergs/s/cm2. This suggests that significant incompleteness is happening in the next lower bin, covering $F_{X}=1-2\times 10^{-10}$ ergs/s/cm2.
@@
We can then look at the distribution of $L_{X}$ in detected outbursts (Fig. 4). (We must keep in mind that the majority of these $L_{X}$ values are likely overestimates, due to the assumed distances being often upper limits, if the observed bursts were Eddington-limited.) The $L_{X}$ distribution looks generally similar to the distribution of fluxes, due to the concentration of transient LMXBs in the Galactic Bulge, but due to the variation in distance, the peak is less sharp. An equivalent plot for BH LMXB systems (Fig. 15 in Tetarenko et al. 2016b [^9] Including a shift down by roughly a factor 3 to move from bolometric to 2-10 keV luminosity.) peaks close to $L_{X}\sim 10^{39}$ erg/s, with a tail extending down to $10^{35}$ erg/s. The difference between BHs and NSs may be attributed to the bias towards longer orbital periods among BHs, which are correlated with higher peak $L_{X}$ (Wu et al., 2010). Black hole X-ray binaries may have longer orbital periods with main sequence donor stars, because unstable mass transfer sets in for mass ratios greater than about 0.8-1.0 for typical stellar structures on the lower main sequence; and at very short periods, even the expected peak luminosities leave black holes in radiatively inefficient states, leading to a strong selection effect against these objects (e.g. Arur & Maccarone 2018). Another selection effect is that verifying a black hole in an LMXB typically requires phase-resolved spectroscopy, which is more difficult in a short-period, thus fainter, system; while verifying a neutron star uses X-ray bursts or pulsations, which aren’t tightly correlated with peak $L_{X}$.
@@
Intriguingly, there is a sharper peak in the $1-2\times 10^{36}$ erg/s bin than appears in the flux histogram. If this peak were due entirely to limited sensitivity below it, we would expect this peak to be less sharp in the $L_{X}$ histogram than the $F_{X}$ histogram. It is possible (but not certain) that the number of outbursts falls off below $10^{36}$ erg/s (”very faint” X-ray binary outbursts). A gap in the (persistent) X-ray luminosity function between $10^{34}$ and $10^{36}$ erg/s has previously been identified in globular clusters (Hertz & Grindlay, 1983; Verbunt & Lewin, 2006; Bahramian et al., 2020b), where the sources above, and just below, this gap consist predominantly of neutron star LMXBs. A dropoff below $L_{X}\sim 10^{36}$ erg/s has also been suggested in the bulge of M31 (Voss & Gilfanov, 2007), in globular clusters in nearby galaxies (Zhang et al., 2011), and possibly (but with inadequate data) for the Milky Way (Sazonov et al., 2006; Bahramian et al., 2021). On the other hand, extensive Swift, XMM-Newton and Chandra monitoring of the Galactic Center identified 28 outbursts from 13 LMXBs, between 2000 and 2014, including 11 between $L_{X}=10^{34}$ erg/s and $10^{35}$ erg/s, which argues strongly against any cutoff (Wijnands et al., 2006; Degenaar et al., 2015). This apparent discrepancy may be reconciled if most persistently-bright systems have $L_{X}>10^{36}$ erg/s, but transient outbursts are often fainter, so that a given snapshot would find a gap, but a list of peak outburst luminosities would not. Following the standard disk-instability model, the only LMXBs that could remain persistently below $10^{36}$ erg/s will be a small range of periods for ultracompact systems in’t Zand et al. (e.g. 2007); van Haaften et al. (e.g. 2013). If this is indeed the case, then more sensitive X-ray monitoring programs (e.g. proposed for STROBE-X, Ray et al. 2019) would find substantially more X-ray outbursts.
@@
Figure 1: Peak fluxes by year from transient neutron star LMXBs. Green, cyan, blue, magenta, and black triangles indicate outbursts by 4U 1608-52, MXB 1730-335, GRS 1747-312, NGC 6440 X-2, and Aquila X-1. Red squares indicate all other transient LMXB outbursts.
@@
Figure 2: Number of total outbursts per year (blue triangles) and outbursts with $F_{X}>3\times 10^{-10}$ ergs/cm2/s per year (red dots), from transient neutron star LMXBs. The effective periods in which the RXTE/ASM, RXTE/PCA bulge scans, INTEGRAL galactic bulge monitoring, Swift/BAT monitor, and MAXI/GSC monitoring were effective are indicated. Since 1996, detections of bright outbursts have been roughly constant in time.
@@ 3.2 Application to the Galactic Center
Hailey et al. (2018) identified a dozen Chandra X-ray sources within a few parsecs of the Galactic Center that have non-thermal X-ray spectra (best fit with a power law with photon index between 1.5 and 2.9), which along with their $L_{X}$ values indicates these are quiescent LMXBs. Hailey et al. (2018) and Mori et al. (2021) suggest that these objects must be black hole LMXBs, arguing that all neutron star LMXBs have short recurrence times, generally less than 5 years, while black hole LMXBs can have longer recurrence times. Maccarone et al. (2022) argued that a number of known neutron star LMXB systems do have longer recurrence times, and that the non-thermal quiescent Galactic Center X-ray sources may be dominated by neutron star LMXBs. Mori et al. (2021) and Mori et al. (2022) argued that most neutron star LMXBs undergo frequent outbursts, based on the history of the known NS transients studied in the Galactic Center. This argument assumes that the confirmed neutron star transients are representative of the underlying population, and not subject to biases; for instance, that the more frequent outbursters are easier to confirm (e.g. they produce more X-ray bursts).
@@
A key question we can ask is what fraction of transient neutron star LMXBs have a recurrence time greater than 10 years, from the data available. We can answer this by checking for gaps of at least 10 years between 1996 and 2022 (the time frame with the best sensitivity). Among the 85 systems we can study, only 24 do not show such a gap, suggesting that 72% could hide for 10 years or longer. Naive direct application of this to the Galactic Centre suggests that for every NS LMXB with a shorter recurrence time (e.g. the 6 known in the Galactic Centre), there are at least three times more systems with longer recurrence periods. This is a lower limit, which does not account for systems that have been quiescent over the past 30+ years, such as the quiescent LMXBs observed in many globular clusters (e.g. Heinke et al., 2003b; Guillot et al., 2011; Bahramian et al., 2015).
@@
This is a very rough estimate, subject to numerous caveats. Most importantly, all-sky monitors have limited sensitivity. We infer that we are missing most outbursts below $L_{X}\sim 2.3\times 10^{36}*(d/(8{\rm kpc}))^{2}$ erg/s; even brighter outbursts for $\sim$10of the year when the Sun is nearby; and some (not easily quantifiable) due to crowding with bright nearby sources. The Swift campaigns of the Galactic Center show that the three verified NS LMXBs there show fainter outbursts as well as bright outbursts; Degenaar et al. (2015) list 4 outbursts with peak $L_{X}>2.3\times 10^{36}$ erg/s, and 11 outbursts with peaks below this. Since 1996, our larger databases give 8 and 22, respectively. There are also nine other transients, showing very-faint X-ray outbursts (peak $L_{X}<10^{36}$ erg/s), that have been detected in the Galactic Centre area since 2000, which have not been confirmed as neutron stars through X-ray bursts, nor show evidence for being black holes (Degenaar et al., 2015). It is generally difficult to attribute X-ray bursts observed near the Galactic Center with RXTE (Galloway et al., 2008), and very faint X-ray binaries may be expected to burst only rarely (Peng et al., 2007; Lin et al., 2019). It seems likely that some (or most) of these very faint X-ray binaries are neutron stars, given that known black hole outbursts tend to be brighter than known neutron star outbursts (Yan & Yu, 2015). However, we have observed few short-period black hole systems, and it is possible that these are relatively faint; cf. Knevitt et al. 2014. These 9 very-faint transients have each had only one or two outbursts within the past 23 years, suggesting relatively long recurrence times ($>$20 years), and that a larger population of such transients may exist there.
@@
For most neutron star transients elsewhere, we cannot constrain their history of fainter outbursts peaking at $L_{X}<2\times 10^{36}$ erg/s. Only for three nearby neutron stars accreting from low-mass companions are we able to rule out outbursts since 1996 above $L_{X}>7\times 10^{34}$ ergs/s; the LMXB Cen X-4 (at 1.3 kpc), and the two transitional millisecond pulsars PSR J1023+0038 (Archibald et al., 2009) and PSR J12270-4859 (Roy et al., 2015), both at 1.4 kpc. It is unclear if these two transitional millisecond pulsars undergo regular, or even any, major accretion outbursts, as IGR 18245-2452 did. These three objects suggest that most neutron stars accreting from low-mass companions remain in a quiescent state for long periods.
@@
Among the other transient neutron star LMXBs with long ($>$10-year) periods between known outbursts, we can identify several more with substantial power-law components in their spectra; XTE J0929-314 (Wijnands et al., 2005b), SAX J1324.5-6313 (Cornelisse et al., 2002a; Bahramian & Heinke, 2015), XB 1733-30 (CX2 in Terzan 1; Cackett et al. 2006; Bozzo et al. 2020, GRO J1744-28 (Doroshenko et al., 2020), and 1RXS J180408.9-342058 (Parikh et al., 2017b).
We note that the XB-NEWS group has conducted frequent optical monitoring of Cen X-4 since 2008 without a proper outburst, although they did observe a slow optical brightening culminating with a failed outburst in late 2020, peaking at $L_{X}\sim 3\times 10^{33}$ erg/s (Baglio et al., 2022); this serves to illustrate that the XB-NEWS optical monitoring (Lewis et al., 2008; Russell et al., 2019) is effective at identifying outbursts. XB-NEWS monitors 50 systems weekly, including a number of other NS LMXB transients detectable in the optical (such as IGR J00291+59, XTE J2123-058, XTE J0929-314, XTE J1814-338, and HETE J1900.1-2455; Lewis et al. 2008). Careful analysis of the XB-NEWS archive may enable determination of how many fainter outbursts we are missing.

### v2

@@ 3.2 General trends
In Fig. 1 we show the peak fluxes from all transient neutron star LMXBs (a few lie outside our bounds), with the most frequent outbursters indicated with different colors. In Fig. 2, we show the number of outbursts per year (omitting Cir X-1’s many outbursts in some years, which can be rather distorting).
@@
From Fig. 2, we see that 1996 marks a dramatic change in detected outburst frequency. A significant peak appears around 2009-2011 in all outbursts. Fig. 2 of Tetarenko et al. (2016b), transient outbursts of BH LMXBs, reveals a similar peak in 2007-2012. However, this peak is not significant in brighter NS LMXB outbursts with $F_{X}>3\times 10^{-10}$ ergs cm-2 s-1, suggesting that fainter outbursts were more detectable in that time frame (e.g. by the RXTE PCA bulge monitoring program; which detected, e.g., the NGC 6440 X-2 outbursts at $1-2\times 10^{-10}$ ergs cm-2 s-1; Heinke et al. 2010).
@@
We then construct a histogram of fluxes from all outbursts since 1996 (Fig. 3, blue triangles). This is sharply peaked at $F_{X}=10^{-9}$ ergs cm-2 s-1, which might suggest that our data becomes highly incomplete below $F_{X}=10^{-9}$ ergs cm-2 s-1. However, if we exclude the four most frequent outbursters (Aquila X-1, 4U 1608-52, MXB 1730-335, and GRS 1747-312), all of which have relatively long orbital periods, high mass transfer rates, and high outburst fluxes, we see a rather different histogram (Fig. 3, red squares), which does not peak until the bin $F_{X}=2-4\times 10^{-10}$ ergs cm-2 s-1. This suggests that significant incompleteness is happening in the next lower bin, covering $F_{X}=1-2\times 10^{-10}$ ergs cm-2 s-1. Referring to Fig. 1, it seems that the majority of monitors have a limit around $F_{X}=2\times 10^{-10}$ ergs cm-2 s-1, so the peak here is likely due to sensitivity limits. Below $F_{X}=1\times 10^{-10}$ ergs cm-2 s-1, most detections use pointed instruments (e.g. Swift/XRT, or RXTE in pointing mode).
@@
We can also look at the distribution of peak $L_{X}$ in detected outbursts (Fig. 4). (We must keep in mind that the majority of these $L_{X}$ values are likely overestimates, due to the assumed distances being often upper limits, if the observed bursts were not actually Eddington-limited as assumed.) The $L_{X}$ distribution looks generally similar to the distribution of fluxes, due to the concentration of transient LMXBs in the Galactic Bulge, but due to the variation in distance, the peak is less sharp. An equivalent plot for BH LMXB systems (Fig. 15 in Tetarenko et al. 2016b [^10] Including a shift down by roughly a factor 3 to move from bolometric to 2-10 keV luminosity.) peaks close to $L_{X}\sim 10^{39}$ erg/s, with a tail extending down to $10^{35}$ erg/s. The tendency for BHs to have higher $L_{X}$ may be attributed to BHs’ bias towards longer orbital periods, which are correlated with higher peak $L_{X}$ (Wu et al., 2010). At very short periods, even the expected peak luminosities leave black holes in radiatively inefficient states, leading to a strong selection effect against low-luminosity black holes (e.g. Arur & Maccarone 2018). Another selection effect is that verifying a black hole in an LMXB typically requires phase-resolved spectroscopy, which is more difficult in a short-period, thus fainter, system; while verifying a neutron star uses X-ray bursts or pulsations, which aren’t tightly correlated with peak $L_{X}$.
@@
Intriguingly, there is a sharper peak in the $1-2\times 10^{36}$ erg/s bin than appears in the flux histogram. If this peak were due entirely to limited sensitivity below it, we would expect this peak to be less sharp in the $L_{X}$ histogram than the $F_{X}$ histogram. It is possible (but not certain) that the number of outbursts falls off below $10^{36}$ erg/s (”very faint” X-ray binary outbursts). A gap in the (persistent) X-ray luminosity function between $10^{34}$ and $10^{36}$ erg/s has previously been identified in globular clusters (Hertz & Grindlay, 1983; Verbunt & Lewin, 2006; Bahramian et al., 2020b), where the sources above, and just below, this gap consist predominantly of neutron star LMXBs. A dropoff below $L_{X}\sim 10^{36}$ erg/s has also been suggested in the bulge of M31 (Voss & Gilfanov, 2007), in globular clusters in nearby galaxies (Zhang et al., 2011), and possibly (but with inadequate data) for the Milky Way (Sazonov et al., 2006; Bahramian et al., 2021). On the other hand, extensive Swift, XMM-Newton and Chandra monitoring of the Galactic Center identified 28 outbursts from 13 LMXBs, between 2000 and 2014, including 11 between $L_{X}=10^{34}$ erg/s and $10^{35}$ erg/s, which argues strongly against any cutoff (Wijnands et al., 2006b; Degenaar et al., 2015). This apparent discrepancy may be reconciled if most persistently-bright systems have $L_{X}>10^{36}$ erg/s, but transient outbursts are often fainter, so that a given snapshot would find a gap, but a list of peak outburst luminosities would not. Following the standard disk-instability model, the only LMXBs that could remain persistently below $10^{36}$ erg/s will be a small range of periods for ultracompact systems (e.g. in’t Zand et al., 2007; van Haaften et al., 2013). If this is indeed the case, then more sensitive X-ray monitoring programs (e.g. proposed for STROBE-X, Ray et al. 2019) would find substantially more X-ray outbursts.
@@
Figure 1: Peak fluxes by year from transient neutron star LMXBs. Green, cyan, blue, magenta, orange, and black triangles indicate outbursts by 4U 1608-52, MXB 1730-335, GRS 1747-312, NGC 6440 X-2, Cir X-1, and Aquila X-1. Red squares indicate all other transient LMXB outbursts. Horizontal colored lines indicate rough sensitivity limits for various X-ray monitoring instruments.
@@
Figure 2: Number of total outbursts per year (blue triangles) and outbursts with $F_{X}>3\times 10^{-10}$ ergs cm-2 s-1 per year (red dots), from transient neutron star LMXBs. The effective periods in which the RXTE/ASM, RXTE/PCA bulge scans, INTEGRAL galactic bulge monitoring, Swift/BAT monitor, and MAXI/GSC monitoring were effective are indicated. Since 1996, detections of bright outbursts have been roughly constant in time.
@@ 3.3 Application to the Galactic Center
Hailey et al. (2018) identified a dozen Chandra X-ray sources within a few parsecs of the Galactic Center that have non-thermal X-ray spectra (best fit with a power law with photon index between 1.5 and 2.9), which along with their $L_{X}$ values indicates these are quiescent LMXBs. Hailey et al. (2018) and Mori et al. (2021) suggest that these objects must be black hole LMXBs, arguing that all neutron star LMXBs have short recurrence times, generally less than 5 years, while black hole LMXBs can have longer recurrence times. Maccarone et al. (2022) argued that a number of known neutron star LMXB systems do have longer recurrence times, and that the non-thermal quiescent Galactic Center X-ray sources may be dominated by neutron star LMXBs. Mori et al. (2021) and Mori et al. (2022) argued that most neutron star LMXBs undergo frequent outbursts, based on the history of the known NS transients studied in the Galactic Center. This argument assumes that the confirmed neutron star transients are representative of the underlying population, and not subject to biases; for instance, that the more frequent outbursters are easier to confirm as neutron stars (e.g. they produce more X-ray bursts).
@@
A key question we can ask is what fraction of transient neutron star LMXBs have a recurrence time greater than 10 years, from the data available. We can answer this by checking for gaps of at least 10 years between 1996 and 2022 (the time frame with the best sensitivity). We identify, for each object, the probability (based on the recorded 1996-2022 outbursts) of not seeing outbursts during a continuous 10-year period. (We allow a continuous time interval to wrap from 2022 to 1996, for this purpose.) For instance, Cen X-4 has a 100% chance of not seeing any outburst in a random 10-year period of this range, while MXB 1659-29 has only a 22% chance of having no outbursts in a 10-year block, and SAX J1808.4-3658 has a 0% chance of seeing no outbursts in a 10-year block. Among 84 systems we study (omitting Terzan 6 X-2, Terzan 5 X-2 and X-3, due to confusion), only 24 do not show any 10-year-long gap, suggesting that the majority have a chance of hiding for 10 years or longer. Randomly selecting a contiguous ten-year period from a random neutron star in our sample gives an average 37% probability of finding no outbursts.
This calculation does not include systems that have been quiescent over the past 30+ years, such as the quiescent LMXBs observed in many globular clusters (e.g. Heinke et al., 2003b; Guillot et al., 2011; Bahramian et al., 2015). As only 15 neutron star LMXBs have been observed to outburst from globular clusters (Bahramian & Degenaar, 2023), but of order 200 quiescent neutron star LMXBs are estimated in Galactic globular clusters (Heinke et al., 2005), this would suggest a neutron star LMXB population over ten times larger than observed. Some evolutionary calculations support the likely existence of such binaries (Meyer-Hofmeister et al., 2024).
The principal caveat is that all-sky monitors have limited sensitivity. We infer that we are missing most outbursts below $L_{X}\sim 2.3\times 10^{36}*(d/(8{\rm kpc}))^{2}$ erg/s; even brighter outbursts for $\sim$10of the year when the Sun is nearby; and some (not easily quantifiable) due to crowding with bright nearby sources. The Swift campaigns on the Galactic Center show that the three closest verified neutron star LMXBs show at least one fainter outburst as well as bright outbursts; Degenaar et al. (2015) list 4 outbursts with peak $L_{X}>2.3\times 10^{36}$ erg/s, and 11 outbursts with peaks below this. Since 1996, our larger databases give 8 and 22, respectively. Mori et al. (2022) argue that the much larger survey frequency of the Galactic Center gives it the only well-constrained outburst histories, particularly regarding very faint X-ray transient behavior.
@@
However, there are also nine other transients, showing very-faint X-ray outbursts (peak $L_{X}<10^{36}$ erg/s), that have been detected in the Galactic Center region since 2000, which have not been confirmed as neutron stars through X-ray bursts, and have also not shown previous evidence for being black holes (Degenaar et al., 2015). These 9 very-faint transients have each had only one or two outbursts within the past 23 years, suggesting relatively long recurrence times ($>$20 years), and that a larger population of such transients may exist there. It is generally difficult to attribute X-ray bursts observed near the Galactic Center with RXTE (Galloway et al., 2008), and very faint X-ray binaries may be expected to burst only rarely (Peng et al., 2007; Lin et al., 2019). The fact that known black hole outbursts tend to be brighter than known neutron star outbursts (Yan & Yu, 2015) suggests that some, or most, of these very faint X-ray binaries are neutron stars. However, we have observed few short-period black hole systems, and it is possible that short-period black hole LMXBs are relatively faint; cf. Knevitt et al. 2014.
One empirical test between neutron star and black hole LMXBs, introduced by Wijnands et al. (2015), compares the photon index of a power-law spectrum fitted to a transient LMXB in the very-faint regime ($L_{X}=10^{34-36}$ erg/s), where black hole LMXBs have significantly harder spectra than most neutron star LMXBs. High extinction weakens the power of this test, but it may still be useful. We have plotted the low-extinction neutron star and black hole systems used by Wijnands et al. (2015) to introduce this test in Fig. 6, along with the measurements published for several high-extinction known neutron star systems (AX J1745.6-2901, GRS 1741.9-2853, SAX J1747.0-2853, KS 1741-293, GRO 1744-28, XMMU J1747161-281048, and XMM J174457-2850.3). All but the last of these systems conform to the trend of high photon indices for neutron star systems with $L_{X}$ between $10^{34-35}$ erg/s. Plotting the very-faint X-ray transients, however, we see that their photon indices align with those of black hole LMXBs. Although this is extremely interesting, we follow Wijnands et al. (2015) in pointing out that a number of neutron star LMXBs also show equally hard spectra in this $L_{X}$ range (e.g. IGR J18245-2452 in M28, Linares et al. 2014; EXO 1745-248, Rivera Sandoval et al. 2018; and XMM J174457-2850.3, Degenaar et al. 2014d). Although some of these are, or may be, transitional millisecond pulsars (Degenaar et al., 2014d), not all such hard-spectrum neutron star LMXBs are, and it seems unlikely that there would be so many transitional millisecond pulsars in active states in such a small region (given that only $\sim$2of quiescent neutron star LMXBs appear as active-state transitional millisecond pulsars, Bahramian et al. 2020b).
Figure 6: Fitted power-law photon indices vs. 0.5-10 keV $L_{X}$, for very faint X-ray transients in the Galactic Center region (blue squares; Degenaar & Wijnands 2010; Degenaar et al. 2012b, 2015), vs. known neutron star LMXBs at similarly high $N_{H}$ (red squares; Degenaar & Wijnands 2010; Degenaar et al. 2012b; Wijnands et al. 2015). Low-$N_{H}$ neutron star LMXBs are plotted as orange crosses, and low-$N_{H}$ black hole LMXBs are plotted as black crosses, both from Wijnands et al. (2015) which introduced the dichotomy. The neutron star LMXB XMM J174457-2850.3 (green triangles), and the other very faint X-ray transients (blue), have unusually hard X-ray spectra, similar to those of black holes.
@@
For most neutron star transients elsewhere, we cannot constrain their history of fainter outbursts peaking at $L_{X}<2\times 10^{36}$ erg/s. Only for three nearby neutron stars accreting from low-mass companions are we able to rule out outbursts since 1996 above $L_{X}>7\times 10^{34}$ ergs s-1; the LMXB Cen X-4 (at 1.3 kpc), and the two transitional millisecond pulsars PSR J1023+0038 (Archibald et al., 2009) and PSR J12270-4859 (Roy et al., 2015), both at 1.4 kpc. It is unclear if these two transitional millisecond pulsars undergo regular, or even any, major accretion outbursts, as IGR 18245-2452 did. These three objects suggest that most neutron stars accreting from low-mass companions remain in a quiescent state for long periods.
@@
Among the other transient neutron star LMXBs with long ($>$10-year) periods between known outbursts, we can identify several more with substantial power-law components in their quiescent spectra; XTE J0929-314 (Wijnands et al., 2005b), SAX J1324.5-6313 (Cornelisse et al., 2002a; Bahramian & Heinke, 2015), XB 1733-30 (CX2 in Terzan 1; Cackett et al. 2006; Bozzo et al. 2020a, GRO J1744-28 (Doroshenko et al., 2020), and 1RXS J180408.9-342058 (Parikh et al., 2017b). XTE J0929-314 and probably XB 1733-30 are ultracompact systems, which would not be dynamically disrupted even in the Galactic Center region.
The XB-NEWS group has conducted frequent optical monitoring of Cen X-4 since 2008 without a proper outburst, although they did observe a slow optical brightening culminating with a failed outburst in late 2020, peaking at $L_{X}\sim 3\times 10^{33}$ erg/s (Baglio et al., 2022); this serves to illustrate that the XB-NEWS optical monitoring (Lewis et al., 2008; Russell et al., 2019b) is effective at identifying outbursts. XB-NEWS monitors 50 systems weekly, including a number of other NS LMXB transients detectable in the optical (such as IGR J00291+59, XTE J2123-058, XTE J0929-314, XTE J1814-338, and HETE J1900.1-2455; Lewis et al. 2008). Careful analysis of the XB-NEWS archive may enable determination of how many fainter outbursts we are missing.